# On estimation and order selection for multivariate extremes via clustering

### v1

@@ 6.2.1 Air Pollution Data
In Figures 6 and 8, following the same manner as in Figure 1, the penalized ASW is plotted against the number of clusters, where different curves correspond to different values of the tuning parameter $t$. With the visual method described in Section 3, we can identify orders as $5$ for the summer data and $3$ for the winter data respectively. These orders are similar to the choices $5$ for the summer data and $4$ for the winter data made in [20] under the guidance of certain elbow plots (see [20, Figure 1]). The authors did not provide a precise explanation of their choices. From the elblow plot in [20, Figure 1], it seems that $k=3$ for the winter data is also plausible. Recall also that here we use the spherical $k$-pc algorithm of [7] while [20] used the spherical $k$-means.
@@
Following the method introduced in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$, we compute the factor coefficient matrix $B$ for the two datasets; see Tables 2 and 2.
Table 1: Estimated $B^{\top}$ for Summer Pollution Data
Factor O3 NO2 NO SO2 PM10
1 0.88 0.22 0.10 0.20 0.24
2 0.20 0.33 0.20 0.90 0.32
3 0.35 0.79 0.30 0.21 0.32
4 0.15 0.16 0.16 0.19 0.80
5 0.21 0.44 0.91 0.25 0.31
Table 2: Estimated $B^{\top}$ for Winter Pollution Data
Factor O3 NO2 NO SO2 PM10
1 0.19 0.98 0.99 0.44 0.98
2 0.07 0.13 0.12 0.89 0.14
3 0.98 0.12 0.10 0.07 0.15
Fig. 6: Penalized ASW Curves for Summer Air Pollution Data
Fig. 7: Squared Cluster Center Coordinates for Summer Air Pollution Data
Fig. 8: Penalized ASW Curves for Winter Air Pollution Data
Fig. 9: Squared Cluster Center Coordinates for Winter Air Pollution Data
@@ 6.2.2 River Discharge Data
Fig. 11: Penalized ASW Curves for River Discharge Data
Fig. 12: Squared Cluster Center Coordinates for River Discharge Data
@@
As in the previous example, Figure 11 presents the penalized ASW curves, from which we found that 6 seems to be an appropriate choice of order. Figure 12 illustrates the squared cluster centers obtained from the $k$-pc algorithm when the order is chosen as $6$. In Table 4, we convert the spectral estimation to the factor matrix $B$ following the method in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$. In addition, for each row of the matrix $B$, we find to which factor index (the same as the cluster index in Figure 12) the largest value (in bold) corresponds. We include these factor indices in the last column of Table 3, which can be viewed roughly as markings of groups of concomitant extremes. These 6 groups are in good accordance with the geographical context: most of the stations located along the same river are found in the same group, with the only exception of the 4 stations along the Mississippi River. The further division of these 4 stations into 2 groups may be easily justified by the large geographical distance between the 2 groups: one group located in MN and the other located in IL.
Factor 1 2 3 4 5 6
SALEM 0.15 0.26 0.11 0.91 0.19 0.18
PORTLAND 0.16 0.27 0.11 0.91 0.19 0.18
HARRISBURG 0.16 0.26 0.12 0.91 0.19 0.17
ST.PAUL 0.88 0.15 0.28 0.10 0.31 0.12
AITKIN 0.91 0.12 0.23 0.13 0.29 0.12
THEBES 0.28 0.15 0.88 0.11 0.31 0.16
CHESTER 0.29 0.15 0.88 0.10 0.30 0.16
BELOW_SPRAGUE 0.22 0.87 0.15 0.25 0.28 0.15
GREEN_ISLAND 0.41 0.26 0.28 0.32 0.69 0.35
FORT_EDWARD 0.31 0.12 0.18 0.16 0.89 0.17
NORTH_CREEK 0.30 0.12 0.17 0.16 0.90 0.18
NEAR_CARLISLE 0.15 0.17 0.15 0.19 0.22 0.92
NEAR_BELL 0.16 0.16 0.14 0.19 0.23 0.92
Table 4: Estimated $B$ for River Discharge Data

### v2

@@ 6.2.1 Air Pollution Data
In Figures 7 and 9, following the same manner as in Figure 1, the penalized ASW is plotted against the number of clusters, where different curves correspond to different values of the tuning parameter $t$. With the visual method described in Section 3, we can identify orders as $5$ (although $4$ seems to be a reasonable choice as well) for the summer data and $3$ for the winter data respectively. These orders are similar to the choices $5$ for the summer data and $4$ for the winter data made in [23] under the guidance of certain elbow plots (see [23, Figure 1]). From the elbow plot in [23, Figure 1], it seems that $k=3$ for the winter data is also plausible. Recall also that here we use the spherical $k$-pc algorithm of [10] while [23] used the spherical $k$-means.
@@
Following the method introduced in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$, we compute the factor coefficient matrix $B$ for the two datasets; see Figures 7 and 9.
Fig. 6: Penalized ASW Curves for Summer Air Pollution Data (top $10\%$ norms).
Fig. 7: Estimated $B^{\top}$ for Summer Pollution Data (top $10\%$ norms).
Fig. 8: Penalized ASW Curves for Winter Air Pollution Data (top $10\%$ norms).
Fig. 9: Estimated $B^{\top}$ for Winter Pollution Data (top $10\%$ norms).
@@ 6.2.2 River Discharge Data
Fig. 11: Penalized ASW Curves for River Discharge Data (top $10\%$ norms)
Fig. 12: Estimated $B^{\top}$ for River Discharge Data (top $10\%$ norms)
@@
As in the previous example, Figure 12 presents the penalized ASW curves, from which we found that 6 seems to be an appropriate choice of order. Figure 12 illustrates the squared cluster centers obtained from the $k$-pc algorithm when the order is chosen as $6$. In Figure 12 , we convert the spectral estimation to the factor matrix $B$ following the method in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$. In addition, for each row of the matrix $B$, we find to which factor index (the same as the cluster index in Figure 12) the largest value (in bold) corresponds. We include these factor indices in the last column of Table 1, which can be viewed roughly as markings of groups of concomitant extremes. The results align well with the expected geographical context: These 6 groups are in good accordance with the geographical context: most of the stations located along the same river are found in the same group, with the only exception of the 4 stations along the Mississippi River. The further division of these 4 stations into 2 groups may be easily justified by the large geographical distance between the 2 groups: one group located in Minnesota (MN) and the other located in Illinois (IL).
In the supplementary material [5], regarding the real data demonstrations presented above, we have included more results with different choices of extremal subsample sizes, or say, top norm percentages. For the air pollution data, the results are relatively stable when the percentage varies. For the river discharge data, when the percentage is decreased to $5\%\sim 1\%$, the penalized ASW criterion starts to favor $k=5$. In this case, the station BELOW SPRAGUE RIVER NEAR CHILOQUIN labeled as group 2 in Table 1 is merged to the concomitant group labeled as 4 formed by SALEM, PORTLAND and HARRISBURG, all of which are geographically close (see Figure 10).
7 Summary and Discussion
Following recent developments in literature, we explore the estimation of multivariate extreme models with a discrete spectral measure by employing spherical clustering techniques. Our main contribution is a method for selecting the order, that is, the number of clusters, which consistently identifies the true number of spectral atoms. Specifically, we introduce an additional penalty term to the well-known simplified average silhouette width, which penalizes small cluster sizes and small dissimilarities between cluster centers. As a by-product, we offer an approach to determining the order of a max-linear factor model. This method is not only straightforward to implement, but also performs effectively in practical applications. We also conduct a large-deviation-type analysis for estimating the discrete spectral measure through clustering methods. This analysis provides insights into the convergence quality of clustering-based estimation for multivariate extremes. In addition, we show how these estimates can be applied to the parameter estimation of heavy-tailed factor models.
At last, we point out several potential future work directions. First, the tuning parameter $t$ for the penalty (22) is chosen with the help of visual inspection. It is of interest to explore the data-driven method to guide the choice of $t$, which may require a refined understanding of the consistency result in Theorem 3.1. In addition, it may be worth exploring alternative clustering assessment criteria to ASW in (21). Some preliminary experiments show that the cross-validation method based on algorithmic instability of [32] might also be a competent criterion in the context of clustering multivariate extremes. It would also be desirable to develop a method for choosing the threshold $l_{n}$ in (14) in the context of clustering multivariate extremes. The recent work [31] might be relevant in this regard. Furthermore, it may be of interest to enhance the large-deviation-type analysis in Section 4 to a finite-sample probabilistic guarantee bound.
Acknowledgment. We thank two anonymous reviewers for their helpful comments that lead to substantial improvements of the article.
Appendix: Supplementary Material
In this supplementary material, we provide results in addition to [4]. In particular, additional simulations studies and real data demonstrations with different choices of extremal subsample thresholds are included.
Appendix A Simulation studies
The results here correspond to Section 6.1 of [4]. We follow the same simulation setup there and the datasets are generated separately independently. The spherical $k$-means and spherical $k$-principal component methods are applied using varying proportions of data with top norms: 1%, 5%, 10%, and 15%, corresponding to effective sample sizes of 100, 500, 1000, and 1500, respectively. The results align with the [4], showing that the proposed penalization has a significant bias correction effect on the selection of the optimal $k$.
A.1 Case: $d=4,k=2$
Fig. 13: Simulation result visualization for the setup $d=4,k=2$ in selection of 1% data
Fig. 14: Simulation result visualization for the setup $d=4,k=2$ in selection of 5% data
Fig. 15: Simulation result visualization for the setup $d=4,k=2$ in selection of 10% data
Fig. 16: Simulation result visualization for the setup $d=4,k=2$ in selection of 15% data
A.2 Case: $d=4,k=6$
Fig. 17: Simulation result visualization for the setup $d=4,k=6$ in selection of 1% data
Fig. 18: Simulation result visualization for the setup $d=4,k=6$ in selection of 5% data
Fig. 19: Simulation result visualization for the setup $d=4,k=6$ in selection of 10% data
Fig. 20: Simulation result visualization for the setup $d=4,k=6$ in selection of 15% data
A.3 Case: $d=6,k=6$:
Fig. 21: Simulation result visualization for the setup $d=6,k=6$ in selection of 1% data
Fig. 22: Simulation result visualization for the setup $d=6,k=6$ in selection of 5% data
Fig. 23: Simulation result visualization for the setup $d=6,k=6$ in selection of 10% data
Fig. 24: Simulation result visualization for the setup $d=6,k=6$ in selection of 15% data
A.4 Case: $d=10,k=6$:
Fig. 25: Simulation result visualization for the setup $d=10,k=6$ in selection of 1% data
Fig. 26: Simulation result visualization for the setup $d=10,k=6$ in selection of 5% data
Fig. 27: Simulation result visualization for the setup $d=10,k=6$ in selection of 10% data
Fig. 28: Simulation result visualization for the setup $d=10,k=6$ in selection of 15% data
Appendix B Real data demonstrations
The results here correspond to Section 6.2 of [4], and we follow exactly the same analysis except for different choices of extremal subsample percentages. The estimated $B^{\top}$ is computed based on the optimal $k$ chosen by inspecting the bending behavior of the corresponding penalized ASW curves.
B.1 Air Pollution Data
The results here correspond to Section 6.2.1 of [4].
Fig. 29: Penalized ASW Curves for Summer Air Pollution Data (10%)
Fig. 30: Estimated $B^{\top}$ (10%)
Fig. 31: Penalized ASW Curves for Summer Air Pollution Data (15%)
Fig. 32: Estimated $B^{\top}$ (15%)
Fig. 33: Penalized ASW Curves for Summer Air Pollution Data (20%)
Fig. 34: Estimated $B^{\top}$ (20%)
Fig. 35: Penalized ASW Curves for Winter Air Pollution Data (10%)
Fig. 36: Estimated $B^{\top}$ (10%)
Fig. 37: Penalized ASW Curves for Winter Air Pollution Data (15%)
Fig. 38: Estimated $B^{\top}$ (15%)
Fig. 39: Penalized ASW Curves for Winter Air Pollution Data (20%)
Fig. 40: Estimated $B^{\top}$ (20%)
B.2 River Discharge Data
The results here correspond to Section 6.2.2 of [4].
Fig. 41: Penalized ASW Curves for River Discharge Data (1%)
Fig. 42: Estimated $B^{\top}$ (1%)
Fig. 43: Penalized ASW Curves for River Discharge Data (5%)
Fig. 44: Estimated $B^{\top}$ (5%)
Fig. 45: Penalized ASW Curves for River Discharge Data (10%)
Fig. 46: Estimated $B^{\top}$ (10%)
Fig. 47: Penalized ASW Curves for River Discharge Data (15%)
Fig. 48: Estimated $B^{\top}$ (15%)
Fig. 49: Penalized ASW Curves for River Discharge Data (20%)
Fig. 50: Estimated $B^{\top}$ (20%)
Appendix C Codes
The R codes that implement the simulation and real data studies can be found at https://github.com/SyuanD/SphCluster.git.